# Retrieval Augmented Generation for Domain-specific Question Answering

### v1

@@
Abstract.
Question answering (QA) has become an important application in the advanced development of large language models. General pre-trained large language models for question-answering are not trained to properly understand the knowledge or terminology for a specific domain, such as finance, healthcare, education, and customer service for a product. To better cater to domain-specific understanding, we build an in-house question-answering system for Adobe products. We propose a novel framework to compile a large question-answer database and develop the approach for retrieval-aware finetuning of a Large Language model. We showcase that fine-tuning the retriever leads to major improvements in the final generation. Our overall approach reduces hallucinations during generation while keeping in context the latest retrieval information for contextual grounding.
Question answering, retrieval, large language model, in-domain understanding.
conference: Workshop on Scientific Document Understanding, February 20–27, 2024, Vancouver, CAN; ; 11footnotetext: Equal Contribution: Authors marked with * contributed equally to this work.
1. Introduction
With the advanced natural language processing technologies, large language models (LLMs) have demonstrated impressive capabilities in a wide range of fields, including chatbots (thoppilan2022lamda, ), language translation (jiao2023chatgpt, ), text summarization (zhang2023benchmarking, ), and personalized robot assistance. Question-answering has become an important application for Large Language Models (LLMs), enhancing the ability of chatbots in diverse domains such as healthcare, finance, education, and customer service (kaddour2023challenges, ).
@@ 1. Introduction
In this work, we develop a new framework to find the answers to the questions that users of Adobe products ask about Adobe products (e.g., ”How can I remove all the highlights in a PDF with Acrobat?”). To address these challenges, our approach builds upon a retrieval-based question-answering system. Our smart retrieval system retrieves the most useful up-to-date relevant content. The retrieval system is trained in a self-supervised manner based on Adobe data and user behavioral data (clicks). We see that fine-tuning the retriever on the domain-specific documents and click data leads to notable improvement in retrieved results nDCG and quality. We then propose a novel approach for retrieval-aware finetuning of a large language model to obtain the question-answering system for Adobe products. We systematically enhance this system to improve its question-answering performance, ensuring users receive timely and accurate information.
@@
  • We proposed a full-fledged, advanced, and production-ready retrieval-based question-answering system, in which the retriever is trained based on user user behavioral data (clicks). We showcase how improving the retriever leads to improvement in the final generation.
@@ 2. Related Work
Large language models have been well developed recently and have shown extraordinary performance in many fields and applications. The models with billions of parameters and trained with huge data can assist humans in simple language understanding or generation tasks. To adapt to a specific task or be aligned with human preference, these models usually need to be finetuned (kaddour2023challenges, ; brown2020language, ). Bakker et al. (bakker2022fine, ) finetune the large language models to align the output of the summarization with humans. In the application in our real life, Singhal et al. evaluate PaLM (chowdhery2022palm, ) with multiple prompting strategies for medical question answering and comprehension. Using a combination of prompting strategies, Flan-PaLM achieves state-of-the-art accuracy on every dataset in the benchmark MultiMedQA. Yu et al. (yu2022legal, ) explore the reason-based prompting mechanism with large language models on the legal question-answering task. Zhou et al. (zhou2023lima, ) propose that strong large language models have already acquired knowledge and capabilities. The model only needs a small amount of data for instruction tuning.
@@
Although the large language model can generate fluent or natural text, it might be unfaithful to the source content, i.e., hallucination (ji2023survey, ; kaddour2023challenges, ; maynez2020faithfulness, ). Parametric knowledge bias is one of the reasons for hallucination in large language models (shuster2021retrieval, ; petroni2019language, ). To mitigate the hallucinations, there are some works that focus on incorporating retrieval-augmented architectures into the systems. Lewis et al. (lewis2020retrieval, ) proposed RAG, a general-purpose fine-tuning recipe for retrieval-augmented generation. RAG combines pre-trained parametric and non-parametric memory for better language generation. Li et al (li2023chatdoctor, ). proposed a medical chat model, ChatDoctor, finetuned on LLaMA using medical domain knowledge. The model is finetuned with the medical data from real-world patient-physician conversations and introductions. With the external knowledge base, the ChatDoctor model can retrieve online and offline medical domain knowledge to answer medical questions on up-to-date medical terms and diseases, mitigating the hallucinations and errors in LLMs.
3. Method
@@ 3.1. Framework Overview
Figure 1. An overview of our proposed framework.
3.2. Retriever
3.2.1. Retriever Training Dataset
@@
Table 1. A sample row from the retriever training dataset.
@@
Table 2. The summary of the dataset.
@@
3.2.2. Retriever Model Training
@@
Figure 2. The training processing for the retriever.
3.3. Retrieval Index Creation and Database
@@
Figure 3. An overall architecture for indexing.
3.4. Preprocessing for Building the Database
@@ 3.4.1. QA Generation Module
We see a significant boost in accuracy in doing few-shot learning compared to zero-shot generation(parnami2022learning, )(song2022comprehensive, )(brown2020language, ). We prepend a couple of sample documents as well as generated QA pairs with each call to our LLM, in order to guide the generation style and quality of the question. In our case, we prefer step-wise responses to questions, as Adobe users tend to prefer discrete steps compared to blocks of text. Table 3 shows a sample QA pair created by the QA generation module.
Table 3. A sample QA pair created by the QA generation module.
@@
3.4.2. Named Entity Removal Module
@@
3.5. Query Augmentation via Product Identification
@@
We utilize a product-intent extraction model(sharma2024semantic, ) that maps all input texts to 1 or more Adobe products. We then pass this information to the retriever, allowing for better relevancy in the retrieved documents.
3.6. LLM prompting
@@
In some cases, some pairs in the list of question-answer pairs may be redundant. Therefore, before the list of question-answer pairs is given to the LLM, our system removes duplicates from the list of question-answer pairs. Specifically, we use it to compute the Levenshtein distance as well as the semantic similarity between each question/title and the list of question-answer pairs. If the Levenstein distance is below some threshold or if the semantic similarity is too high, we compute the semantic similarity between their two answers. If the two answers are too semantically similar, we remove one of them based on the source credibility (Helpx ¿ Community ¿ YouTube ¿ LLM-generated QAs) and answer length.
3.7. LLM Finetuning
@@
$y=\text{LLM}_{\theta}(d^{+},d^{-},q)$. $y$ is the generated outcome, and it is compared to $a$ to update the LLM. The following shows our template to generate the training samples.
@@
4. Experiments
4.1. Retriever Evaluation
@@
For our retriever evaluation, we create evaluation datasets per product, verified by human annotators to ensure quality. For each evaluation, we choose queries Adobe users have asked before (sampling on query length and query frequency). We made sure the model has not seen the queries previously in training.
  1. (1)
    Acrobat Query-Title Match - Create a 300 query to Helpx document title dataset. We sampled both short and long queries when constructing this dataset.
  2. (2)
    Photoshop Query-Title Match - 250 queries to helpx document title dataset.
  3. (3)
    Lightroom Query-Title Match - 100 queries to helpx document title dataset.
  4. (4)
    Adobe Express Query-Title Match - 250 queries to helpx document title dataset.
While computing the nDCG metrics for our finetuned model, we consider all documents in our document set rather than just documents in our evaluation batch to make the task challenging. We compare our retriever to several open and closed sourced options available in industry.
We see significant improvement in retrieved accuracy with our finetuned model, with a smaller model size as shown in table 4.
Table 4. Retriever performance evaluated on nDCG
Model Acrobat (k=4) Photoshop (k=4) Lightroom (k=4) Express (k=4) Model Size
Ours (Finetuned Retriever) 0.7281 0.6921 0.7249 0.8221 109m
BGE-large-en-v1.5 (2023) 0.7031 0.5680 0.6894 0.8143 335m
UAE-Large-V1 (2023) 0.7161 0.5601 0.6743 0.8153 335m
DiffCSE (roberta-base-sts, 2022) 0.2850 0.2571 0.3547 0.3162 124m
SimCSE (sup-roberta-large, 2021) 0.5705 0.3755 0.5786 0.6384 355m
Clip (ViT-L/14) 0.0970 0.0349 0.3459 0.1619 427m
OpenClip (ViT-H-14) 0.3794 0.2867 0.4699 0.6813 986m
T5 (3b) 0.1887 0.1454 0.2959 0.1872 1.24b
SFR-Embedding-Mistral (2024) 0.1623 0.1545 0.2726 0.1377 7b
mxbai-embed-large-v1 (2024) 0.7150 0.5526 0.6783 0.8175 335m
@@
  2. (2)
    Ability to comprehend shorter sentences. Our user behavior shows that users often type sentences ¡ 15 words to get the data they need. Our model outperforms the baseline on shorter sentences.
@@
Table 5. Summary of the retrieval index database.
@@
4.2. Query Augmentation via Product Identification
@@
Table 6. The generated response with and without product identification for the retriever.
@@
To create a blank PDF, select File ¿ New From Template, open the Blank Templates folder, select a template, and select New. Illustrator creates a new document based on the selected blank template.
@@
1. Open Acrobat and choose “Tools” ¿ “Create PDF”.
@@
4.3. Quantitative Evaluation
In addition to doing component specific evaluation, we also evaluate the full system both quantitatively and qualitatively. For our quantitative evaluation, we focus on Adobe Acrobat help questions. We utilize 137 user-asked questions and utilize Acrobat product-experts to generate gold responses for each questions. The questions include short common questions such as ”How to remove password from the PDF”, as well as longer questions such as ”I have a large size pdf which I a unbale to share. how do I compress it and share this compressed pdf with others”.
For the evaluation questions, we generate responses from all candidates and then utilize GPT4(openai2023gpt4, ) to evaluate the similarity of the candidate-generated answer with the gold human-annotated answer on a scale of 0-1. For GPT-4, we set ’n = 20, temperature = 1, top p = 1’ to sample 20 times to estimate the token probabilities(liu2023geval, ). Finally, we take the mean of all scores for a row. The relevance scores are showcased in table 7.
We see that improving the retriever does affect overall quality of the generations as well. For in-domain tasks, we recommend finetuning the retriever to enhance overall generation performance. We also see poor performance of the finetuned generator on the test set. We ascribe this due to a gap between LLM-generated qa pairs (trainset) and human-generated ground-truth writing style.
Table 7. Generation Relevance on Acrobat test set.
Model Relevance Score (0-1)
Finetuned Retriever + Finetuned Generator 0.5150
Finetuned Retriever + GPT 4 0.7242
GPT4 0.1705
SFR-Embedding-Mistral + GPT4 0.7022
sentence-transformers/multi-qa-MiniLM-L6-cos-v1 + GPT4 0.5625
mxbai-embed-large-v1 + GPT4 0.6617
4.4. Qualitative Evaluation - Comparison of the generated answers
@@
    As of my last knowledge update in September 2021, there is no widely recognized software or product called ”Adobe Firefly” from Adobe Inc. Adobe is known for its extensive range of software products, including Adobe Photoshop, Adobe Illustrator, Adobe Premiere Pro, Adobe Acrobat, and many others, which are commonly used for design, video editing, document management, and various creative and productivity tasks. […]

### v2

# AAAI 24’, Workshop on Scientific Document Understanding, February 20–27, 2024, Vancouver, CAN

@@
Abstract
Question answering has become an important application in the advanced development of large language models. The general pre-trained large language models for question-answering are not trained to properly understand the knowledge or terminology for a specific domain, such as finance, healthcare, education, and customer service for a product. As these systems usually have no access to the latest information and cannot adjust to customized features the product expects, we build an in-house question-answering system for our Adobe products. We propose a novel framework to compile a large question-answer database and develop the approach for retrieval-aware finetuning of a Large Language model. Our overall approach reduces hallucinations during generation while keeping in context the latest retrieval information for contextual grounding.
keywords:
[…]
1 Introduction
With the advanced natural language processing technologies, large language models (LLMs) have demonstrated impressive capabilities in a wide range of fields, including chatbots [thoppilan2022lamda], language translation [jiao2023chatgpt], text summarization [zhang2023benchmarking], and personalized robot assistance. Question-answering has become an important application for Large Language Models (LLMs), enhancing the ability of chatbots in diverse domains such as healthcare, finance, education, and customer service [kaddour2023challenges].
@@ 1 Introduction
In this work, we develop a new framework to find the answers to the questions that users of Adobe products ask about Adobe products (e.g., "How can I remove all the highlights in a PDF with Acrobat?"). To address these challenges, our approach builds upon a retrieval-based question-answering system. Our smart retrieval system retrieves the most useful up-to-date relevant content. The retrieval system is trained in a self-supervised manner based on Adobe data and user behavioral data (clicks). We then propose a novel approach for retrieval-aware finetuning of a large language model to obtain the question-answering system for Adobe products. We systematically enhance this system to improve its question-answering performance, ensuring users receive timely and accurate information.
@@
  • We proposed a full-fledged, advanced, and production-ready retrieval-based question-answering system, in which the retriever is trained based on user user behavioral data (clicks).
@@ 2 Related Work
Large language models have been well developed recently and have shown extraordinary performance in many fields and applications. The models with billions of parameters and trained with huge data can assist humans in simple language understanding or generation tasks. To adapt to a specific task or be aligned with human preference, these models usually need to be finetuned [kaddour2023challenges, brown2020language]. Bakker et al. [bakker2022fine] finetune the large language models to align the output of the summarization with humans. In the application in our real life, Singhal et al. evaluate PaLM [chowdhery2022palm] with multiple prompting strategies for medical question answering and comprehension. Using a combination of prompting strategies, Flan-PaLM achieves state-of-the-art accuracy on every dataset in the benchmark MultiMedQA. Yu et al. [yu2022legal] explore the reason-based prompting mechanism with large language models on the legal question-answering task. Zhou et al. [zhou2023lima] propose that strong large language models have already acquired knowledge and capabilities. The model only needs a small amount of data for instruction tuning.
@@
Although the large language model can generate fluent or natural text, it might be unfaithful to the source content, i.e., hallucination [ji2023survey, kaddour2023challenges, maynez2020faithfulness]. Parametric knowledge bias is one of the reasons for hallucination in large language models [shuster2021retrieval, petroni2019language]. To mitigate the hallucinations, there are some works that focus on incorporating retrieval-augmented architectures into the systems. Lewis et al. [lewis2020retrieval] proposed RAG, a general-purpose fine-tuning recipe for retrieval-augmented generation. RAG combines pre-trained parametric and non-parametric memory for better language generation. Li et al [li2023chatdoctor]. proposed a medical chat model, ChatDoctor, finetuned on LLaMA using medical domain knowledge. The model is finetuned with the medical data from real-world patient-physician conversations and introductions. With the external knowledge base, the ChatDoctor model can retrieve online and offline medical domain knowledge to answer medical questions on up-to-date medical terms and diseases, mitigating the hallucinations and errors in LLMs.
3 Method
@@ 3.1 An overview of our framework
Figure 1: An overview of our proposed framework.
3.2 Retriever
3.2.1 Retriever Training Dataset
@@
Table 1: A sample row from the retriever training dataset.
@@
Table 2: The summary of the dataset.
@@
3.2.2 Retriever Model Training
@@
Figure 2: The training processing for the retriever.
3.3 Retrieval Index Creation and Database
@@
Figure 3: An overall architecture for indexing.
3.4 Preprocessing for Building the Database
@@ 3.4.1 QA Generation Module
We see a significant boost in accuracy in doing few-shot learning compared to zero-shot generation[parnami2022learning][song2022comprehensive][brown2020language]. We prepend a couple of sample documents as well as generated QA pairs with each call to our LLM, in order to guide the generation style and quality of the question. In our case, we prefer step-wise responses to questions, as Adobe users tend to prefer discrete steps compared to blocks of text. Table 3 shows a sample QA pair created by the QA generation module.
Table 3: A sample QA pair created by the QA generation module.
@@
3.4.2 Named Entity Removal Module
@@
3.5 Query Augmentation via Product Identification
@@
We utilize a product-intent extraction model that maps all input texts to 1 or more Adobe products. We then pass this information to the retriever, allowing for better relevancy in the retrieved documents.
3.6 LLM prompting
@@
In some cases, some pairs in the list of question-answer pairs may be redundant. Therefore, before the list of question-answer pairs is given to the LLM, our system removes duplicates from the list of question-answer pairs. Specifically, we use it to compute the Levenshtein distance as well as the semantic similarity between each question/title and the list of question-answer pairs. If the Levenstein distance is below some threshold or if the semantic similarity is too high, we compute the semantic similarity between their two answers. If the two answers are too semantically similar, we remove one of them based on the source credibility (Helpx > Community > YouTube > LLM-generated QAs) and answer length.
3.7 LLM Finetuning
@@
$y=\text{LLM}_{θ}(d^{+},d^{-},q)$. $y$ is the generated outcome, and it is compared to $a$ to update the LLM. The following shows our template to generate the training samples.
@@
4 Experiments
4.1 Retriever Evaluation
@@
We reserved  7% of rows from our overall user behavioral data to use for evaluation. While computing the nDCG metrics for our finetuned model, we consider all documents in our document set rather than just documents in our evaluation batch to make the task challenging. Our evaluation data contained 47936 rows. The nDCG for both the Pretrained Model (MPnet) and our finetuned model are shown in table 4
Table 4: Retriever performance evaluated on nDCG for the pretrained MPnet and our finetuned model.
# of Evaluation Rows nDCG Pretrained Model (MPnet) nDCG Finetuned Model (Ours)
47,936 0.23 0.391
@@
  2. 2.
    Ability to comprehend shorter sentences. Our user behavior shows that users often type sentences < 15 words to get the data they need. Our model outperforms the baseline on shorter sentences.
@@
Table 5: Summary of the retrieval index database.
@@
4.2 Query Augmentation via Product Identification
@@
Table 6: The generated response with and without product identification for the retriever.
@@
To create a blank PDF, select File > New From Template, open the Blank Templates folder, select a template, and select New. Illustrator creates a new document based on the selected blank template.
@@
1. Open Acrobat and choose “Tools” > “Create PDF”.
@@
4.3 Quantitative Evaluation
In addition to doing component specific evaluation, we also evaluate the full system both quantitatively and qualitatively. For our quantitative evaluation, we focus on Adobe Acrobat help questions. We utilize 137 user-asked questions and utilize Acrobat product-experts to generate gold responses for each questions. The questions include short common questions such as "How to remove password from the PDF", as well as longer questions such as "I have a large size pdf which I a unbale to share. how do I compress it and share this compressed pdf with others".
For the evaluation questions, we generate responses from all candidates and then utilize GPT4[openai2023gpt4] to evaluate the similarity of the candidate-generated answer with the gold human-annotated answer on a scale of 1-5. For GPT-4, we set ’n = 20, temperature = 1, top p = 1’ to sample 20 times to estimate the token probabilities[liu2023geval]. Finally, we take the mean of all scores for a row. The relevance scores are showcased in table 7.
Table 7: The relevance scores for all generation candidate models on Acrobat test set.
Model Relevance Score (1-5)
GPT4 3.39
Finetuned Retriever + GPT4 4.24
4.4 Qualitative Evaluation - Comparison of the generated answers
@@
    As of my last knowledge update in September 2021, there is no widely recognized software or product called "Adobe Firefly" from Adobe Inc. Adobe is known for its extensive range of software products, including Adobe Photoshop, Adobe Illustrator, Adobe Premiere Pro, Adobe Acrobat, and many others, which are commonly used for design, video editing, document management, and various creative and productivity tasks. […]
@@
Example 2
ChatGPT might fail to generate the correct answer for some questions related to the process of how to use Adobe products, while our system can generate the step-by-step answer.
  • Question: How can I customize my birthday poster using Adobe Express?
  • Answer from ChatGPT:
    As of my last knowledge update in September 2021, there isn’t a specific application called “Adobe Express.” However, Adobe offers a range of creative applications, and it’s possible that a new one has been released since then. Adobe Spark Post is a popular tool for creating posters, and it’s quite user-friendly. Here’s a general guide on how to customize a birthday poster using Adobe Spark Post or a similar Adobe application: […]
  • Answer from our system:
    Steps:
    1. Open Adobe Express for free on your desktop or mobile device.
    2. Choose from the available birthday poster templates or create from a blank canvas.
    3. Upload your own photos or add stock images and designs from the libraries.
    4. Brand your poster by uploading your logo and using fonts and colors that match your aesthetic.
    5. Share your poster by hitting the publish button and instantly downloading it to your device. Print it out or share it online.
Example 3
In some cases, ChatGPT knows the answer. However, it is rather verbose and it adds some information that is useless to users. In this example, ChatGPT added useless information to Acrobat DC users such as “Please note that Adobe Acrobat DC is a paid software, and you may need a subscription or license to access certain features. Ad”. By contrast, the answer of our system is concise, correct, and does not add any useless information:
  • Question: How can remove one page in a PDF in Acrobat DC?
  • Answer from ChatGPT:
    However, it is rather verbose and all it adds some information that is useless to Acrobat DC users such as “Please note that Adobe Acrobat DC is a paid software, and you may need a subscription or license to access certain features. Ad”
  • Answer from our system:
    Steps:
    1. Choose the “Organize Pages” tool.
    2. Select the page thumbnail of any page or pages you want to delete, then click the “Delete” icon to remove the page or pages from the file.
    3. After clicking the “Delete” icon, click “OK” to apply changes to remove the pages.
    4. Save the file.
@@
Example 5
This example shows that OpenAI’s ChatGPT fails to give the proper answer to the question because of missing the fact that the user wants to trim the video before adding it to the timeline.
Our system is better suited for this more expert-level knowledge and provides the correct answer.
  • Question: How do I trim a video before adding it in the timeline in Premiere Pro?
  • Answer from ChatGPT:
    ChatGPT missed the fact that the user wants to trim the video before adding it in the timeline.
  • Answer from our system: Steps:
    1. Double-click on any of the media in the project panel to preload it into the source monitor.
    2. Play or scrub the media.
    3. Mark an endpoint by hitting the I key or clicking the icon.
    4. Drag your video to where you’d like to have an outpoint, either hitting the mark out icon or pressing O.